OpenStudy (boxman61):

solve quadratic equations- Solve: 2x^2+1=0

OpenStudy (anonymous):

complete the square or use the quadratic equation.

OpenStudy (anonymous):

quadratic formula

OpenStudy (boxman61):

this is what i've done so far \[2x ^{2}+1=0\]\[2x ^{2}=-1\]\[x ^{2}=-1/2\]\[x=\sqrt{-1/2}\] answer is \[\pm \sqrt{2}/2i\]

OpenStudy (anonymous):

\[x^{2} = -1/2\] \[x = \pm (\sqrt{2}/2) i \]

OpenStudy (boxman61):

thats the right answer glgan1, what step am i missing?

OpenStudy (anonymous):